# Skewness-dependent Moments of Pion GPD from Nonlocal Quark-Bilinear Correlators

### v1

@@ IV.2 Extraction of moments at zero-skewness
With the fit range and $t$-dependence parameterizations established above, we begin with the zero-skewness case, which serves as the cleanest benchmark for our analysis. Combined fits to the renormalized matrix elements $\widetilde{\mathcal{H}}^{R}$ are performed using data from three boosted momenta, each covering six values of the momentum transfer t-t. The GFFs are parameterized using both the monopole form of Eq. (25) and the $z$-expansion of Eq. (26). Both parameterizations provide good descriptions of the lattice data, yielding $\chi^{2}/\mathrm{d.o.f.}=1.17$ (monopole) and $0.91$ ($z$-expansion). A detailed comparison of the two approaches is given in Appendix C. Unlike the VMD-inspired monopole form, the $z$-expansion is more flexible. In particular, it produces more conservative uncertainty estimates, as reflected in its broader error bands. For these reasons, we adopt the $z$-expansion in the subsequent analysis.
Figure 6: Demonstration of the combined fit to the original lattice data at $\xi=0$. The left and right panels display the results with $t=0~\mathrm{GeV}^{2}$ and $t\approx 0.2~\mathrm{GeV}^{2}$, respectively, each including results for three different momenta.
@@
Our results for the lowest three odd moments are shown in Fig. 7, which compares the fixed-order (NNLO) determinations with the NNLL-resummed analyses. The fixed-order results are shown as solid-filled bands. For the NNLL-resummed results, the hatched bands contain both the statistical and systematic uncertainties from varying $\kappa$ between 1 and 2. Specifically, we use bootstrap resampling to construct confidence intervals and quantify the uncertainty. For each of the three scale parameters $\kappa=\{1,\,\sqrt{2},\,2\}$, we generate 200 bootstrap samples and compute the standard deviations. The final uncertainty bands are then defined by the envelope of the upper and lower bounds derived from all three $\kappa$ values, thus combining the statistical uncertainty with the systematic uncertainty arising from $\kappa$ variation. In particular, $A_{1,0}$ (left panel) corresponds to the pion electromagnetic form factor. Previous lattice studies [144] have demonstrated that both the pion charge radius and form factor are sensitive to the pion mass used in lattice simulations. To ensure a meaningful comparison, we therefore restrict our $A_{1,0}$ results to those of Ref. [144] at $m_{\pi}=300~\mathrm{MeV}$, obtaining reasonable agreement. The remaining discrepancies with experimental determinations [21, 153, 152, 151, 149, 150] are attributable to the relatively heavy pion mass used in our lattice calculations.
@@ IV.3 Extraction of moments with skewness dependence
$\Delta\widetilde{\mathcal{H}}(\lambda,\xi,t)$ $=\xi^{2}\left[-2\,\lambda^{2}\,A_{3,2}(t)\,c_{3,0}+\frac{\lambda^{4}}{6}\Big(A_{5,2}(t)\,c_{5,0}+\xi^{2}\big(A_{3,2}(t)\,c_{5,2}+4\,A_{5,4}(t)\,c_{5,0}\big)\Big)\right].$ (28)
@@
Figure 8: Comparison between the subtracted lattice data and the fitted $\xi$-dependent contribution $\Delta\widetilde{\mathcal{H}}^{\text{fit}}(\lambda,\xi,t)$ at different values of skewness $\xi$ and similar t-t. For clarity, the left panel includes $\xi=-0.11$ and $\xi=-0.14$, while the right panel includes $\xi=-0.25$ and $\xi=-0.33$. In each panel, the two datasets have similar t-t values. Markers denote the subtracted lattice data, and solid bands show the fit results.
While the subtraction test in Eq. (28) serves only as an intermediate diagnostic, the central physical information is provided by the fitted GFFs and reconstructed Mellin moments, which are presented in Figs. 912. As for NLL–resummation, the coupled evolution equations for the matching coefficients, Eq. (13), are solved numerically using a Runge–Kutta algorithm, with the two-loop anomalous-dimension matrix and two-loop $\beta$ function included. The fitted results are shown in Fig. 7, where the bands include the statistical uncertainty and the scale-variation uncertainty estimated by varying $\kappa$ between 1 and 2.
@@
Figure 11: Results for $A_{5,0}(t)$, $A_{5,2}(t)$, and $A_{5,4}(t)$ (are shown as blue bands in each subpanel) obtained from combined fits to the zero- and nonzero-skewness datasets. The gray bands in the left panel show the $A_{5,0}(t)$ results obtained from the $\xi=0$ dataset, with the solid band corresponding to the NLL-resummed extraction and the hatched band corresponding to the NNLL-resummed one, for comparison with the combined-fit determination.
@@ ACKNOWLEDGMENTS
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract No. DE-SC0012704, Contract No. DE-AC02-06CH11357, and within the frameworks of Scientific Discovery through Advanced Computing (SciDAC) award Fundamental Nuclear Physics at the Exascale and Beyond and the Topical Collaboration in Nuclear Theory Quark Gluon Tomography (QGT) with Award DE-SC0023646. QS is supported by Kent State University through index 201452.
@@ Appendix A Two-loop anomalous dimension and analytic evolution in the zero-skewness limit
$\hat{C}^{\overline{\mathrm{MS}}}(z^{2}\mu^{2})$ $=\hat{C}_{00}+\alpha_{s}\left(\hat{C}_{11}L+\hat{C}_{10}\right)+\alpha_{s}^{2}\left(\hat{C}_{22}L^{2}+\hat{C}_{21}L+\hat{C}_{20}\right)+\cdots,$ (31)
where all coefficients $\hat{C}_{nm}$ with $\{n,m=0,1,2\}$ are matrices. Specifically, $\hat{C}_{n2}$, $\hat{C}_{n1}$, and $\hat{C}_{n0}$ denote the matrices collecting the coefficients of the double-logarithmic, single-logarithmic, and finite terms, respectively, with $\hat{C}_{00}$ being the identity matrix for the case of moments. The RG equation for $\hat{C}^{\overline{\mathrm{MS}}}(z^{2}\mu^{2})$ reads:
@@
$C_{n+1,k}\!\left(z^{2}\mu^{2}\right)\;\equiv\;i^{\,k}\binom{n}{k}\int_{0}^{1}\!d\alpha\,\alpha^{\,n-k}\,\left.\frac{d^{\,k}}{d(\xi\lambda)^{k}}K\!\left(\alpha,\xi\lambda,z^{2}\mu^{2}\right)\right|_{\lambda=0}.$ (38)
@@
$C_{n+1,k}\!\left(z^{2}\mu^{2}\right)\;\equiv\;i^{\,k}\binom{n}{k}\int_{0}^{1}\!d\alpha d\beta\,(1-\alpha-\beta)^{\,n-k}\left.\frac{d^{\,k}}{d(\xi\lambda)^{k}}\left(K(\alpha,\beta,z^{2}\mu^{2})e^{-i(\alpha-\beta)\xi\lambda}\right)\right|_{\lambda=0}.$ (39)
@@
$C_{1,0}$ $=1+\alpha_{s}\left(0.318\,L+0.743\right)+\alpha_{s}^{2}\left(0.165\,L^{2}+1.034\,L+1.028\right),$
$C_{2,0}$ $=1+\alpha_{s}\left(0.601\,L-0.106\right)+\alpha_{s}^{2}\left(0.396\,L^{2}+0.372\,L+1.341\right),$
$C_{3,0}$ $=1+\alpha_{s}\left(0.760\,L-0.654\right)+\alpha_{s}^{2}\left(0.561\,L^{2}-0.348\,L+2.464\right),$
$C_{4,0}$ $=1+\alpha_{s}\left(0.874\,L-1.088\right)+\alpha_{s}^{2}\left(0.694\,L^{2}-1.042\,L+3.838\right),$
$C_{5,0}$ $=1+\alpha_{s}\left(0.962\,L-1.456\right)+\alpha_{s}^{2}\left(0.807\,L^{2}-1.701\,L+5.326\right),$
$C_{3,2}$ $=\alpha_{s}\left(0.195-0.088\,L\right)+\alpha_{s}^{2}\left(-0.079\,L^{2}+0.195\,L+0.209\right),$
$C_{4,2}$ $=\alpha_{s}\left(0.279-0.117\,L\right)+\alpha_{s}^{2}\left(-0.128\,L^{2}+0.434\,L-0.014\right),$
$C_{5,2}$ $=\alpha_{s}\left(0.354-0.134\,L\right)+\alpha_{s}^{2}\left(-0.164\,L^{2}+0.651\,L-0.313\right),$
$C_{5,4}$ $=\alpha_{s}\left(0.053-0.028\,L\right)+\alpha_{s}^{2}\left(-0.022\,L^{2}+0.042\,L+0.107\right).$ (40)
@@
In the zero-skewness case, the perturbative expression and anomalous dimension reduce to the standard DGLAP limit, for which analytic evolution kernels are available. At LL accuracy, the evolution reads [^3] To make the expressions more compact, we will use $a_{s}(\mu)\equiv\alpha_{s}(\mu)/(4\pi)$ in place of $\alpha_{s}$.
@@
$C_{i,0}^{\rm RG}(a_{s}(\mu),z^{2}\mu^{2})$ $=C_{i,0}(a_{s}(\mu_{0}),z^{2}\mu_{0}^{2})\left(\frac{a_{s}(\mu_{0})}{a_{s}(\mu)}\right)^{\frac{\Upsilon_{0}^{ii}}{\beta_{0}}}\exp\!\left\{\frac{\beta_{0}\Upsilon_{1}^{ii}-\beta_{1}\Upsilon_{0}^{ii}}{\beta_{0}\beta_{1}}\ln\!\frac{a_{s}(\mu_{0})\beta_{1}+\beta_{0}}{a_{s}(\mu)\beta_{1}+\beta_{0}}\right\}.$ (42)
@@
$C_{i,0}^{\rm RG}(a_{s}(\mu),z^{2}\mu^{2})$ $=C_{i,0}(a_{s}(\mu_{0}),z^{2}\mu_{0}^{2})\,\left(\frac{a_{s}(\mu_{0})}{a_{s}(\mu)}\right)^{\frac{\Upsilon_{0}^{ii}}{\beta_{0}}}\exp\!\Bigg\{\frac{\beta_{0}\Upsilon_{1}^{ii}-\beta_{1}\Upsilon_{0}^{ii}}{\beta_{0}\beta_{1}}\ln\!\frac{\beta_{1}+\beta_{0}\,a_{s}(\mu_{0})}{\beta_{1}+\beta_{0}\,a_{s}(\mu)}$ (43)
$\qquad\qquad+\frac{1}{2\beta_{0}}\int_{a_{s}(\mu)}^{a_{s}(\mu_{0})}\!da_{s}\,\bigg[\frac{\Upsilon_{2}^{ii}}{a_{s}}-\frac{\beta_{2}\Upsilon_{0}^{ii}+\beta_{1}\Upsilon_{1}^{ii}}{-\beta_{0}a_{s}^{2}-\beta_{1}a_{s}^{3}-\beta_{2}a_{s}^{4}}\bigg]\Bigg\}.$
@@ Appendix B Extraction of the Lorentz-invariant amplitudes
Figure 13: Extracted LI amplitudes for two representative nonzero-skewness cases. Left panel: $P_{z}=1.937$ GeV, t=0.864-t=0.864 GeV2, right panel: $P_{z}=1.937$ GeV, t=2.748-t=2.748 GeV2.

### v2

@@ IV.2 Extraction of moments at zero-skewness
With the fit range and $t$-dependence parameterizations established above, we begin with the zero-skewness case, which serves as the cleanest benchmark for our analysis. Combined fits to the renormalized matrix elements $\widetilde{\mathcal{H}}^{R}$ are performed using data from three boosted momenta, each covering six values of the momentum transfer t-t. The GFFs are parameterized using both the monopole form of Eq. (25) and the $z$-expansion of Eq. (26). Both parameterizations provide good descriptions of the lattice data, yielding $\chi^{2}/\mathrm{d.o.f.}=1.17$ (degrees of freedom) (monopole) and $0.91$ ($z$-expansion). A detailed comparison of the two approaches is given in Appendix C. Unlike the VMD-inspired monopole form, the $z$-expansion is more flexible. In particular, it produces more conservative uncertainty estimates, as reflected in its broader error bands. For these reasons, we adopt the $z$-expansion in the subsequent analysis.
Figure 6: Demonstration of the combined fit to the original lattice data at $\xi=0$. The upper and lower panels display the results with $t=0~\mathrm{GeV}^{2}$ and $t\approx 0.2~\mathrm{GeV}^{2}$, respectively, each including results for three different momenta.
@@
Our results for the lowest three odd moments are shown in Fig. 7, which compares the fixed-order (NNLO) determinations with the NNLL-resummed analyses. The fixed-order results are shown as solid-filled bands. For the NNLL-resummed results, the hatched bands contain both the statistical and systematic uncertainties from varying $\kappa$ between 1 and 2. Specifically, we use bootstrap resampling to construct confidence intervals and quantify the uncertainty. For each of the three scale parameters $\kappa=\{1,\,\sqrt{2},\,2\}$, we generate 200 bootstrap samples to estimate the statistical uncertainties, which is sufficient to yield stable error estimates for our dataset. The final uncertainty bands are then defined by the envelope of the upper and lower bounds derived from all three $\kappa$ values, thus combining the statistical uncertainty with the systematic uncertainty arising from $\kappa$ variation. In particular, $A_{1,0}$ (left panel) corresponds to the pion electromagnetic form factor. Previous lattice studies [147] have demonstrated that both the pion charge radius and form factor are sensitive to the pion mass used in lattice simulations. To ensure a meaningful comparison, we therefore restrict our $A_{1,0}$ results to those of Ref. [147] at $m_{\pi}=300~\mathrm{MeV}$, obtaining reasonable agreement. The remaining discrepancies with experimental determinations [22, 156, 155, 154, 152, 153] are attributable to the relatively heavy pion mass used in our lattice calculations.
@@ IV.3 Extraction of moments with skewness dependence
$\Delta\widetilde{\mathcal{H}}(\lambda,\xi,t)=\xi^{2}$ $\bigg[-2\,\lambda^{2}\,A_{3,2}(t)\,c_{3,0}+\frac{\lambda^{4}}{6}\Big(A_{5,2}(t)\,c_{5,0}$
$+\xi^{2}\big(A_{3,2}(t)\,c_{5,2}+4\,A_{5,4}(t)\,c_{5,0}\big)\Big)\bigg].$ (28)
@@
Figure 8: Comparison between the subtracted lattice data and the fitted $\xi$-dependent contribution $\Delta\widetilde{\mathcal{H}}^{\text{fit}}(\lambda,\xi,t)$ at different values of skewness $\xi$ and similar t-t. For clarity, the upper panel includes $\xi=-0.11$ and $\xi=-0.14$, while the lower panel includes $\xi=-0.25$ and $\xi=-0.33$. In each panel, the two datasets have similar t-t values. Markers denote the subtracted lattice data, and solid bands show the fit results.
While the subtraction test in Eq. (IV.3) serves only as an intermediate diagnostic, the central physical information is provided by the fitted GFFs and reconstructed Mellin moments, which are presented in Figs. 912. As for NLL–resummation, the coupled evolution equations for the matching coefficients, Eq. (13), are solved numerically using a Runge–Kutta algorithm, with the two-loop anomalous-dimension matrix and two-loop $\beta$ function included. The fitted results are shown in Fig. 7, where the bands include the statistical uncertainty and the scale-variation uncertainty estimated by varying $\kappa$ between 1 and 2.
@@
Figure 11: Results for $A_{5,0}(t)$, $A_{5,2}(t)$, and $A_{5,4}(t)$ (shown as blue bands in each subpanel) obtained from combined fits to the zero- and nonzero-skewness datasets. The gray bands in the left panel show the $A_{5,0}(t)$ results obtained from the $\xi=0$ dataset, with the solid band corresponding to the NLL-resummed extraction and the hatched band corresponding to the NNLL-resummed one, for comparison with the combined-fit determination.
@@ ACKNOWLEDGMENTS
We thank Yushan Su and Yuxun Guo for the valuable discussions. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract No. DE-SC0012704, Contract No. DE-AC02-06CH11357, and within the frameworks of Scientific Discovery through Advanced Computing (SciDAC) award Fundamental Nuclear Physics at the Exascale and Beyond and the Topical Collaboration in Nuclear Theory Quark Gluon Tomography (QGT) with Award DE-SC0023646. QS is supported by Kent State University through index 201452.
@@ Appendix A Two-loop anomalous dimension and analytic evolution in the zero-skewness limit
$\hat{C}^{\overline{\mathrm{MS}}}(z^{2}\mu^{2})=$ $\hat{C}_{00}+\alpha_{s}\left(\hat{C}_{11}L_{z}+\hat{C}_{10}\right)$
$+\alpha_{s}^{2}\left(\hat{C}_{22}L_{z}^{2}+\hat{C}_{21}L_{z}+\hat{C}_{20}\right)+\cdots,$ (31)
where $L_{z}=\ln(\mu^{2}z^{2})$, and all coefficients $\hat{C}_{nm}$ with $\{n,m=0,1,2\}$ are matrices. Specifically, $\hat{C}_{n2}$, $\hat{C}_{n1}$, and $\hat{C}_{n0}$ denote the matrices collecting the coefficients of the double-logarithmic, single-logarithmic, and finite terms, respectively, with $\hat{C}_{00}$ being the identity matrix for the case of moments. The RG equation for $\hat{C}^{\overline{\mathrm{MS}}}(z^{2}\mu^{2})$ reads:
@@
$C_{n+1,k}$ $(z^{2}\mu^{2})$
$\,=\,$ $i^{k}\!\binom{n}{k}\!\int_{0}^{1}\!d\alpha\,\alpha^{\,n-k}\!\left.\frac{d^{k}}{d(\xi\lambda)^{k}}K(\alpha,\xi\lambda,z^{2}\mu^{2})\right|_{\lambda=0}.$ (38)
@@
$C_{n+1,k}\!\left(z^{2}\mu^{2}\right)\!=$ $i^{\,k}\binom{n}{k}\int_{0}^{1}\!d\alpha d\beta\,(1-\alpha-\beta)^{\,n-k}$
$\times\frac{d^{\,k}}{d(\xi\lambda)^{k}}\left(K(\alpha,\beta,z^{2}\mu^{2})e^{-i(\alpha-\beta)\xi\lambda}\right)\bigg|_{\lambda=0}.$ (39)
@@
$C_{1,0}$ $=1+\alpha_{s}\left(0.318\,L_{z}+0.743\right)+\alpha_{s}^{2}\left(0.165\,L_{z}^{2}+1.034\,L_{z}+1.028\right),$
$C_{2,0}$ $=1+\alpha_{s}\left(0.601\,L_{z}-0.106\right)+\alpha_{s}^{2}\left(0.396\,L_{z}^{2}+0.372\,L_{z}+1.341\right),$
$C_{3,0}$ $=1+\alpha_{s}\left(0.760\,L_{z}-0.654\right)+\alpha_{s}^{2}\left(0.561\,L_{z}^{2}-0.348\,L_{z}+2.464\right),$
$C_{4,0}$ $=1+\alpha_{s}\left(0.874\,L_{z}-1.088\right)+\alpha_{s}^{2}\left(0.694\,L_{z}^{2}-1.042\,L_{z}+3.838\right),$
$C_{5,0}$ $=1+\alpha_{s}\left(0.962\,L_{z}-1.456\right)+\alpha_{s}^{2}\left(0.807\,L_{z}^{2}-1.701\,L_{z}+5.326\right),$
$C_{3,2}$ $=\alpha_{s}\left(0.195-0.088\,L_{z}\right)+\alpha_{s}^{2}\left(-0.079\,L_{z}^{2}+0.195\,L_{z}+0.209\right),$
$C_{4,2}$ $=\alpha_{s}\left(0.279-0.117\,L_{z}\right)+\alpha_{s}^{2}\left(-0.128\,L_{z}^{2}+0.434\,L_{z}-0.014\right),$
$C_{5,2}$ $=\alpha_{s}\left(0.354-0.134\,L_{z}\right)+\alpha_{s}^{2}\left(-0.164\,L_{z}^{2}+0.651\,L_{z}-0.313\right),$
$C_{5,4}$ $=\alpha_{s}\left(0.053-0.028\,L_{z}\right)+\alpha_{s}^{2}\left(-0.022\,L_{z}^{2}+0.042\,L_{z}+0.107\right).$ (40)
@@
In the zero-skewness case, the perturbative expression and anomalous dimension reduce to the standard DGLAP limit, for which analytic evolution kernels are available. At LL accuracy, the evolution reads, where we use $a_{s}(\mu)\!\equiv\!\alpha_{s}(\mu)/(4\pi)$ in place of $\alpha_{s}$ for compactness:
@@
$C_{i,0}^{\rm RG}(a_{s}(\mu),$ $z^{2}\mu^{2})=C_{i,0}(a_{s}(\mu_{0}),z^{2}\mu_{0}^{2})\left(\frac{a_{s}(\mu_{0})}{a_{s}(\mu)}\right)^{\frac{\Upsilon_{0}^{ii}}{\beta_{0}}}$
$\times\exp\!\left\{\frac{\beta_{0}\Upsilon_{1}^{ii}-\beta_{1}\Upsilon_{0}^{ii}}{\beta_{0}\beta_{1}}\ln\!\frac{a_{s}(\mu_{0})\beta_{1}+\beta_{0}}{a_{s}(\mu)\beta_{1}+\beta_{0}}\right\}.$
@@
$C_{i,0}^{\rm RG}$ $(a_{s}(\mu),z^{2}\mu^{2})=C_{i,0}(a_{s}(\mu_{0}),z^{2}\mu_{0}^{2})\,\left(\frac{a_{s}(\mu_{0})}{a_{s}(\mu)}\right)^{\frac{\Upsilon_{0}^{ii}}{\beta_{0}}}$
$\exp\!\Bigg\{\frac{\beta_{0}\Upsilon_{1}^{ii}-\beta_{1}\Upsilon_{0}^{ii}}{\beta_{0}\beta_{1}}\ln\!\frac{\beta_{1}+\beta_{0}\,a_{s}(\mu_{0})}{\beta_{1}+\beta_{0}\,a_{s}(\mu)}+\frac{1}{2\beta_{0}}$
$\quad\times\int_{a_{s}(\mu)}^{a_{s}(\mu_{0})}\!da_{s}\bigg[\frac{\Upsilon_{2}^{ii}}{a_{s}}\!-\!\frac{\beta_{2}\Upsilon_{0}^{ii}+\beta_{1}\Upsilon_{1}^{ii}}{-\beta_{0}a_{s}^{2}-\beta_{1}a_{s}^{3}-\beta_{2}a_{s}^{4}}\bigg]\Bigg\}.$
@@ Appendix B Extraction of the Lorentz-invariant amplitudes
Figure 13: Extracted LI amplitudes for two representative nonzero-skewness cases. Upper panel: $P_{z}=1.937$ GeV, t=0.864-t=0.864 GeV2, lower panel: $P_{z}=1.937$ GeV, t=2.748-t=2.748 GeV2.